imbadatmath:

Perform the operation. (3x^2-7x)-(4x^2+6x)

supie:

OK

supie:

=-X^2-13X

why:

Let's simplify step-by-step. 3x2−7x−(4x2+6x) Distribute the Negative Sign: =3x2−7x+−1(4x2+6x) =3x2+−7x+−1(4x2)+−1(6x) =3x2+−7x+−4x2+−6x Combine Like Terms: =3x2+−7x+−4x2+−6x =(3x2+−4x2)+(−7x+−6x) =−x2+−13x Answer: =−x2−13x

supie:

yeah

imbadatmath:

it sayed it was wrong

supie:

oh ok

imbadatmath:

then it swithched to this

imbadatmath:

Perform the operation. (4x^2+6x-5)-(-7x^2-4x-2)

supie:

is that the right 1?

imbadatmath:

it sayed it was wrong and instead of giving me the answer it gave me a different question

imbadatmath:

so its new

imbadatmath:

do you know